# Automatic Mapping of AutomationML Files to Ontologies for Graph Queries and Validation

### v1

@@ 4 AutomationML Ontology
As described in Section 3.2, a new AutomationML ontology was designed that aims to to fulfill the Requirements outlined in Section 3.1. The section first gives an overview of the CAEX vocabulary that could be reused (R2), before describing the concepts that had to be added to fulfill Requirements R3 and R4.
4.1 Reused entities from previous AutomationML Ontology
The AutomationML Ontology from Kovalenko et al. [3] already offered strong coverage of the basic CAEX concepts. Because of this, most of the common concepts could be reused in the new ontology (s. Table 3). Many additional concepts are present in the ontology, but needlessly differ from the AutomationML terminology. These have been renamed to align them with Requirement R1.
AutomationML Concept | Ontology Class |
CAEXFile | :CAEXFile |
InstanceHierarchy | :InstanceHierarchy |
SystemUnitClassLib | :SystemUnitClassLib |
RoleClassLib | :RoleClassLib |
InterfaceClassLib | :InterfaceClassLib |
InternalElement | :InternalElement |
SystemUnitClass | :SystemUnitClass |
RoleClass | :RoleClass |
InterfaceClass | :InterfaceClass |
ExternalInterface | :ExternalInterface |
Attribute | :Attribute |
ExternalReference | :ExternalReference |
InternalLink | :InternalLink |
AutomationML Concept | Object Property |
InternalStructure | :hasPart, :hasInterface, :hasAttribute |
InternalLink | :hasRefPartnerSideA, :hasRefPartnerSideB |
RoleRequirements | :hasRoleRequirement |
SupportedRoleClass | :hasSupportedRoleClass |
4.2 Additional CAEX Vocabulary
In part due to continued development of the AutomationML-Standard, some parts of the CAEX vocabulary had to be added to fulfill Requirement R2. This mostly refers to concepts needed to define the semantics of attributes. By combining these new concepts with the reused concepts from Table 3, full coverage of the CAEX vocabulary according to R2 can be achieved. Since the new concepts cover the changes introduced by the new AutomationML version, the resulting ontology also covers Requirement R4.
AttributeTypeLibrary and AttributeTypes
The recent AutomationML-Standard IEC 62714-2[14] introduced both AttributeTypes and AttributeTypeLibraries. These can be used to give Attributes of e.g. InternalElements a consistent and reusable semantic description. They are structured in the same way as e.g. the InterfaceClasses and InterfaceClassLibraries. Therefore, in order to allow the modeling of AttributeTypes in the ontology, two classes were added to the AutomationML-Ontology: AttributeType, which allows the modeling of a template that can be reused for Attributes and AttributeTypeLib, which contains and defines AttributeTypes. These classes are linked to the remaining concepts of the ontology through newly defined object properties: The object property hasRefAttributeType is a direct translation of the CAEX-Tag RefAttributeType. In AutomationML it serves two purposes: First, it links an Attribute to the AttributeType it is based on. Second, it can also link an AttributeType to the AttributeType it is derived from.
Constraints
In AutomationML, the values of attributes can be restricted using three different types of constraints:
A NominalScaledType-constraint restricts the values of an attribute to a set of allowed values, while an OrdinalScaledType-constraint restricts the values to a range specified by a minimal and a maximal attribute type. Additionally, an UnknownType-constraint exists that can be used for any other type of constraint.
In order to allow the use of these constraints in the AutomationML-Ontology, the classes NominalScaledConstraint, OrdinalScaledConstraint and UnknownTypeConstraint were added to the ontology, alongside data properties to define their constraints.
MappingObjects
In AutomationML, MappingObjects map the attributes or interfaces of an InternalElement to the corresponding attributes or interfaces defined in RoleClasses or SystemUnitClasses. If the names of the Attributes or Interfaces match, this is done implicitly, otherwise an explicit MappingObject is created. This ensures that the data and functionalities defined in the InternalElement are correctly associated with the corresponding elements in the RoleClass/SystemUnitClass. This allows for clear and consistent interpretation of the roles and properties within the AutomationML model.
4.3 Context dependent Vocabulary
On top of the CAEX vocabulary, AutomationML uses regular CAEX concepts which can have different meanings depending of their context. To end users, these concepts are only ever shown in their contextual meaning (e.g. as a Facet instead of an InternalElement). In order to increase usability for the end user, these relationships should be made explicit in the ontology in accordance with Requirement R3.
MirrorObjects
In AutomationML, a MirrorObject is used to reference and replicate elements from one part of the model into another without duplicating the actual data. It creates a ”mirror” of an existing object, such as an InternalElement, SystemUnitClass, or other elements, allowing the same object to be used in multiple locations within the model.
The MirrorObject references the original object by its ID, ensuring that any changes to the original are automatically reflected in all mirrored instances. This promotes consistency and reduces redundancy within the AutomationML model by reusing existing data structures instead of copying them. MirrorObjects are not part of the CAEX vocabulary, but can instead be interpreted based on their context. In order to be able to explicitly model Mirror Object relationships, the object properties hasMirrorObject and hasMasterObject have been added to the ontology. These properties connect InternalElements or Interfaces with their respective Mirror/Master-Objects.
Facet
A Facet is a mechanism used to group related attributes, interfaces, and sub-elements within a single object. It allows for the modular and reusable definition of properties and behaviors, which can then be applied to various elements in the model, enhancing consistency and simplifying the management of complex systems. Facets help in organizing and structuring the data more efficiently by encapsulating related aspects of a model element. While Facets are not part of the CAEX vocabulary, they are commonly used in AutomationML to create domain specific views of a component and are therefore explicitly included as a class in the ontology.
Group
In AutomationML, a Group is a mechanism used to logically organize and manage multiple elements, such as InternalElements or SystemUnitClasses, into a collective structure. Groups allow for the aggregation of related elements, facilitating easier handling, reuse, and understanding of complex models within a project. Like Facets, they are not part of the CAEX-vocabulary, but are explicitly included in the ontology as a class.
@@ 4.4 Design Practices and Reuse
The AutomationML ontology is published under W3ID, which is a permanent identifier, that is administered by the W3C Permanent Identifier Community Group. It is permanently available using the ontology base IRI w3id.org/hsu-aut/AutomationML. Administration and maintenance of the ontology is supported by the Institute of Automation Technology at Helmut-Schmidt-University. Both the ontology itself as well as its classes and relationships are annotated with metadata documenting their creation process and correct usage.
4.5 Ontology Summary
The AutomationML-Ontology contains a total of 22 classes, 8 ObjectProperties and 5 Data Properties. Their terminology remains very close to the terms used in AutomationML (R1). Together, these classes and properties allow the full coverage of both the CAEX-Vocabulary according to Requirement R2 as well as the AutomationML-specific vocabulary according to Requirement R3. By updating the terms to cover all newly added terms from the most recent specification, it allows full coverage of the current version of AutomationML (R4).
@@ 5 Mapping AutomationML to OWL
While Section 4 assured that the AutomationML-Ontology provides a comprehensive vocabulary to cover the entirety of the AutomationML-content, the following Section provides:
@@
An alignment of AutomationML vocabulary with OWL-specific vocabulary (e.g. rdf:type etc.),
@@ 5.4 Mapping of Class Hierarchies
In the SystemUnitClass-, RoleClass-, InterfaceClass-, and AttributeTypeLibraries, this inheritance mechanism results in a class hierarchy using rdfs:subClassOf according to Equations 6 and 7.
@@ 5.10 Inheritance of Subcomponents
In AutomationML, the internal structure of a class is inherited from a parent class to the child class. In the AutomationML model, this relationship is not explicitly modeled, but can instead be inferred through the inheritance relationship. It is additionally possible to override inherited objects.
For the mapping to RDF, this means that an internal class structure is inherited from the parent class if the inherited structure was not overridden (s. Equation 14).
@@ 6 Mapping Example
To illustrate the mapping rules from the previous section, a Four-Tank Mixing System will be used as a running example. Further models are available on Github.
@@ 6.2 Creation of IRIs
Listing 6.2 and Figure 4 show an example of this logic.
While the InternalElement B201 is assigned an IRI from its ID, both the Attribute Length as well as the AttributeType Length use their paths as an IRI.
{listing}[ht]
Example of an InternalElement with an Attribute ”Length”.
{Verbatim}[commandchars=
¡InstanceHierarchy Name=”MyIH”¿
¡InternalElement Name=”B201”
ID=”44806a23-d2bd-45d2-8344”¿
¡Attribute Name=”Length”
RefAttributeType=”MyAtLib/Length”¿
¡/Attribute¿
¡/InternalElement¿
¡/InstanceHierarchy¿
@@ 6.3 Creation of Classes and Instances
The following section provides an example of the creation of classes and instances. Listing 6.3 shows an example of an InternalElement, that has all major elements that AutomationML uses: A SystemUnitClass, an Attribute, an ExternalInterface as well as two RoleClasses - one as a SupportedRoleClass and one as a RoleRequirement.
[ht]
InternalElement with all basic AutomationML elements.
{Verbatim}[commandchars=
¡InternalElement Name=”B201”
ID=”44806a23-d2bd-45d2-8344”
RefBaseSystemUnitPath=”MySucLib/Vessel”¿
¡Attribute Name=”Length”
RefAttributeType=”MyAtLib/Length”¿
¡/Attribute¿
¡ExternalInterface Name=”Input”
ID=”ce45d7b3-169d-4be8-9746”
RefBaseClassPath=”MyIcLib/Port”¿
¡/ExternalInterface¿
¡SupportedRoleClass
RefRoleClassPath=”MyRcLib/Vessel” /¿
¡RoleRequirements
RefBaseRoleClassPath=”MyRcLib/DosingTank” /¿
¡/RoleRequirements¿
¡/InternalElement¿
Figure 5 shows the corresponding relationship between OWL instances and classes from Listing 6.3: For example, the InternalElement B201 is:
@@ 6.4 Mapping of Class Hierarchies
The example in Listing 6.4 and Figure 6 shows how the AttributeType Length from the previous example is mapped and embedded into an OWL class hierarchy. For AttributeTypes, the rdfs:subClassOf-relationship can be inferred from the RefAttributeType in AutomationML.
[ht]
Excerpt from the AttributeLibrary.
{Verbatim}[commandchars=
¡AttributeTypeLib Name=”MyAtLib”¿
¡AttributeType
Name=”Dimensions”¿
¡AttributeType Name=”Length”
RefAttributeType=”MyAtLib/Dimensions” /¿
¡/AttributeType¿
¡/AttributeTypeLib¿
@@ 6.5 Mapping of InternalLinks
Listing 6.5 shows an example of two InternalElements, whose interfaces are connected by an InternalLink. Figure 7 shows both mechanisms of linking their Interfaces. On one hand, they are connected in a way that very closely follows the AutomationML-vocabulary with an instance of the class InternalLink that connects the Interfaces through the Object Property :hasRefPartnerSideA/B. On the other hand, they are connected directly through the object property :isLinked. This makes traversing the connections between InternalElements (e.g. for queries) more intuitive, while preserving the terminology most familiar to AutomationML practitioners.
[ht]
Example of an InternalLink that connects a Pipe with a Tank.
{Verbatim}[commandchars=
¡InstanceHierarchy Name=”MyIH”¿
¡InternalElement Name=”Pipe”
ID=”a20e3eac-9ec0-41f1-852a”¿
¡ExternalInterface Name=”Output”
ID=”6eea7a40-43fd-4aee-a1d3-”¿
¡/ExternalInterface¿
¡InternalLink Name=”Pipe˙B201”
RefPartnerSideA=”6eea7a40-43fd-4aee-a1d3”
RefPartnerSideB=”ce45d7b3-169d-4be8-9746”/¿
¡/InternalElement¿
¡InternalElement Name=”B201”
ID=”44806a23-d2bd-45d2-8344”¿
¡ExternalInterface Name=”Input”
ID=”ce45d7b3-169d-4be8-9746”¿
¡/ExternalInterface¿
¡/InstanceHierarchy¿
7 Usage
@@ 7.1 Complex Graph Queries
However, many industrial applications could significantly benefit from the possibility to query the full spectrum of information provided by the AutomationML file.
@@
{Verbatim}[commandchars=
//InternalElement[RoleRequirements[
@RefBaseRoleClassPath=’MyRcLib/Vessel’]]
@@
{Verbatim}[commandchars=
for $ie in //InternalElement
where $ie/RoleRequirements/
@RefBaseRoleClassPath = ”MyRcLib/Vessel”
return $ie
SPARQL:
{Verbatim}[commandchars=
SELECT ?ie
WHERE –
?ie rdf:type aml:InternalElement .
?ie aml:hasRoleRequirement aml:MyRcLib/Vessel.˝
The queries regarding the first scenario were easily achievable in all three languages (s. Listing 8).
[ht]
xQuery for Scenario 2 and SPARQL Query for Scenario 3.
xQuery:
{Verbatim}[commandchars=
let $roleClasses :=
//RoleClass[@RefBaseClassPath=’PlantEquipment’]
for $ie in //InternalElement
where some $rc in $roleClasses satisfies
$ie/RoleRequirements/@RefBaseRoleClassPath =
string-join($rc/ancestor-or-self::*/@Name,’/’)
return $ie
SPARQL:
{Verbatim}[commandchars=
SELECT ?ie
WHERE –
?ie rdf:type aml:InternalElement .
?ie aml:hasRoleRequirement/(rdfs:subClassOf)*
aml:PlantEquipment .˝
Both xPath and xQuery operate directly on the XML-tree, and have therefore no mechanisms to handle ReferencePaths. xQuery offers some mechanisms to circumvent this limitation: Technically, the second scenario could still be solved using xQuery. The approach includes reconstructing the reference path of all RoleClasses, iterating over them, and individually comparing them to the desired value (s. Listing 8). However, this approach is very limited in scope, e.g. capturing the open-ended transitive closure statement from Query 3 in this way is not possible. Using SPARQL, due to the support of Property Paths, all three queries can be answered in a much more compact and arguably more intuitive form (s. Listing 8).
@@
A SPARQL query that solves the second scenario can be found in Listing 9.
[ht]
SPARQL query, that creates a material flow graph while considering valve states.
{Verbatim}[commandchars=
construct–?s aml:flows ?o.˝ where –
?s a aml:InternalElement.
?o a aml:InternalElement.
?s aml:hasInterface ?i1.
?o aml:hasInterface ?i2.
?i1 aml:isLinkedTo ?i2.
OPTIONAL –
?s aml:hasAttribute ?vs.
?vs aml:hasRefAttributeType aml:ValveState.
?vs aml:hasAttributeValue ?state.
Filter(!bound(?state) —— ?state = ’true’).
@@ 7.2 Data Validation
In this use case, the validation concerns the correct usage of elements that were imported from user defined libraries. These libraries are commonly accompanied by textual specifications that define the proper usage of the imported elements. As those specifications lack a formal or machine interpretable definition, their correct usage needs to be manually checked by domain experts. For this use case, the Application Recommendation Automation Project Configuration (AR APC) was considered. It is a framework that describes guidelines for the configuration of automation projects.[31]
@@
{Verbatim}[commandchars=
¡rule context=”caex:InternalElement[
caex:RoleRequirements[@RefBaseRoleClassPath=
’ArApcRcLib/SensorPort’]]”¿
¡assert test=”
every $link in caex:InternalLink satisfies
(//caex:ExternalInterface[
@ID=$link/@RefPartnerSideA]/@RefBaseClassPath=
’ArApcIcLib/SensorPortInterface’ and
//caex:ExternalInterface[
@ID=$link/@RefPartnerSideB]/@RefBaseClassPath=
’ArApcIcLib/SensorPortInterface’)”¿
¡/assert¿
¡/rule¿
SHACL:
{Verbatim}[commandchars=
:SPIConnectionShape
a sh:NodeShape ;
sh:targetClass aml:SensorPortInterface;
sh:property [
sh:path aml:isLinkedTo;
sh:class aml:SensorPortInterface;] .

### v2

@@ 4 AutomationML Ontology
To address the requirements outlined in Section 3.1, a new ontology has been developed. This section introduces the AutomationML ontology by outlining its intended scope, presenting its main contents, and explaining the underlying design principles that guided its development.
4.1 Ontology Scope
AutomationML is structured around a clear separation between the metamodel and the model level.
The metamodel is defined by the CAEX schema (IEC 62424) together with AutomationML-specific constraints. It specifies the abstract language of AutomationML: which elements exist (InternalElement, SystemUnitClass, RoleClass, InterfaceClass, Attribute), how they may be related, and which properties they can carry. In ontological terms, this corresponds to the T-Box, i.e., the definition of the vocabulary and its structural rules. The core ontology introduced in this work captures precisely this metamodel.
On the model level, AutomationML distinguishes between two types of content:
Libraries (e.g., RoleClassLib, SystemUnitClassLib, InterfaceLib, AttributeTypeLib) define reusable class hierarchies. They extend the metamodel vocabulary with domain- or project-specific categories. In ontology terms, these libraries represent T-Box extensions, since they define new classes and relations, e.g., MixingTank as a subclass of Tank.
Instances populate these libraries with concrete elements that describe a system. For example, a project may instantiate a particular MixingTank as part of a specific plant. These correspond to A-Box assertions, since they describe individual objects and their relationships.
Because libraries and instances are typically user-defined and vary across projects, it is not feasible to provide a predefined ontology that covers them comprehensively. Instead, the ontology developed here focuses on the core vocabulary of AutomationML (metamodel, T-Box). Section 5 then specifies the rules for extending this core ontology with project-specific libraries (T-Box) and populating it with project instances (A-Box).
@@ 4.1 Ontology Scope
As described in Section 3.2, a new AutomationML ontology was designed that aims to to fulfill the Requirements outlined in Section 3.1. The section first gives an overview of the CAEX vocabulary that could be reused (R2.), before describing the concepts that had to be added to fulfill Requirements R4. and R3..
4.2 Ontology Summary
As discussed in Section 3.2, the AutomationML ontology developed by Kovalenko et al. [3] already provides substantial coverage of the core CAEX terms. This ontology was therefore chosen as the starting point for the development of the present AutomationML ontology.
Several modifications and extensions were applied to better align the ontology with current requirements:
Terminology alignment: The terminology was adjusted to closely match the official AutomationML vocabulary (R1.).
Version updates: New concepts introduced in the latest AutomationML version, such as AttributeTypes, were incorporated (R4.).
Documentation and annotations: Additional descriptions and annotations were added to improve clarity and usability.
The ontology follows the AutomationML metamodel and retains its original terminology. For instance, a RoleClass continues to reference its parent class using the hasRefBaseClass object property. The corresponding ontology concept, rdfs:subClassOf, is applied as a separate object property, in accordance with the mapping rules defined in Section 5.
A graphical overview of the classes and object properties in the AutomationML ontology is shown in Figure 4.
@@ 4.3 Design Practices and Reuse
The AutomationML ontology is published under W3ID, which is a permanent identifier, that is administered by the W3C Permanent Identifier Community Group. The ontology is permanently available using the ontology base IRI https://w3id.org/hsu-aut/AutomationML. Administration and maintenance of the ontology is supported by the Institute of Automation Technology at Helmut-Schmidt-University. Both the ontology itself as well as its classes and relationships are annotated with metadata documenting their creation process and correct usage.
@@ 5 Mapping AutomationML to OWL
While Section 4 described the ontology that covers the AutomationML metamodel, the following section defines how the metamodel is extended by a class hierarchy. With that aim, the following Section provides:
@@
An alignment of AutomationML vocabulary with OWL-specific vocabulary (e.g. rdf:type and rdfs:subClassOf),
@@ 5.4 Mapping of Class Hierarchies
In the SystemUnitClass-, RoleClass-, InterfaceClass-, and AttributeTypeLibraries, this inheritance mechanism results in a class hierarchy using rdfs:sub-ClassOf according to Equations 6 and 7.
@@ 5.10 Inheritance of Subcomponents
In AutomationML, the internal structure of a class is inherited from a parent class to the child class. In the AutomationML model, this relationship is not explicitly modeled, but can instead be inferred through the inheritance relationship. It is additionally possible to override inherited objects. For the mapping to RDF, this means that an internal class structure is inherited from the parent class if the inherited structure was not overridden (s. Equation 14).
@@ 6 Mapping Example
To illustrate the mapping rules from the previous section, a Four-Tank Mixing System will be used as a running example. Further models and test cases are available on Github.
@@ 6.2 Creation of IRIs
Listing 6 and Figure 7 show an example of this logic. While the InternalElement B201 is assigned an IRI from its ID, both the Attribute Length as well as the AttributeType Length use their paths as an IRI.
<InstanceHierarchy Name="MyIH"> <InternalElement Name="B201" ID="44806a23-d2bd-45d2-8344"> <Attribute Name="Length" RefAttributeType="MyAtLib/Length"> </Attribute> </InternalElement></InstanceHierarchy>
@@ 6.3 Creation of Classes and Instances
The following section provides an example of the creation of classes and instances. Listing 8 shows an example of an InternalElement, that has all major elements that AutomationML uses: A SystemUnitClass, an Attribute, an ExternalInterface as well as two RoleClasses - one as a SupportedRoleClass and one as a RoleRequirement.
<InternalElement Name="B201" ID="44806a23-d2bd-45d2-8344" RefBaseSystemUnitPath="MySucLib/Vessel"> <Attribute Name="Length" RefAttributeType="MyAtLib/Length"> </Attribute> <ExternalInterface Name="Input" ID="ce45d7b3-169d-4be8-9746" RefBaseClassPath="MyIcLib/Port"> </ExternalInterface> <SupportedRoleClass RefRoleClassPath="MyRcLib/Vessel" /> <RoleRequirements RefBaseRoleClassPath="MyRcLib/DosingTank" /> </RoleRequirements></InternalElement>
Figure 9 shows the corresponding relationship between OWL instances and classes from Listing 8: For example, the InternalElement B201 is:
@@ 6.4 Mapping of Class Hierarchies
The example in Listing 10 and Figure 11 shows how the AttributeType Length from the previous example is mapped and embedded into an OWL class hierarchy. For AttributeTypes, the rdfs:subClassOf-relationship can be inferred from the RefAttributeType in AutomationML.
<AttributeTypeLib Name="MyAtLib"> <AttributeType Name="Dimensions"> <AttributeType Name="Length" RefAttributeType="MyAtLib/Dimensions" /> </AttributeType></AttributeTypeLib>
@@ 6.5 Mapping of InternalLinks
Listing 12 shows an example of two InternalElements, whose interfaces are connected by an InternalLink. Figure 13 shows both mechanisms of linking their Interfaces. On one hand, they are connected in a way that very closely follows the AutomationML-vocabulary with an instance of the class InternalLink that connects the Interfaces through the Object Property :hasRefPartnerSideA/B. On the other hand, they are connected directly through the object property :isLinked. This makes traversing the connections between InternalElements (e.g. for queries) more intuitive, while preserving the terminology most familiar to AutomationML practitioners.
<InstanceHierarchy Name="MyIH"> <InternalElement Name="Pipe" ID="a20e3eac-9ec0-41f1-852a"> <ExternalInterface Name="Output" ID="6eea7a40-43fd-4aee-a1d3-"> </ExternalInterface> <InternalLink Name="Pipe_B201" RefPartnerSideA="6eea7a40-43fd-4aee-a1d3" RefPartnerSideB="ce45d7b3-169d-4be8-9746"/> </InternalElement> <InternalElement Name="B201" ID="44806a23-d2bd-45d2-8344"> <ExternalInterface Name="Input" ID="ce45d7b3-169d-4be8-9746"> </ExternalInterface></InstanceHierarchy>
7 Application
@@ 7.1 Complex Graph Queries
However, many industrial applications could significantly benefit from the possibility to query the full spectrum of information provided by the Automation-ML-file.
@@
//InternalElement[RoleRequirements[ @RefBaseRoleClassPath=’MyRcLib/Vessel’]]xQuery:
for $ie in //InternalElement where $ie/RoleRequirements/ @RefBaseRoleClassPath = "MyRcLib/Vessel" return $ieSPARQL:
SELECT ?ieWHERE { ?ie rdf:type aml:InternalElement . ?ie aml:hasRoleRequirement aml:MyRcLib/Vessel.}
The queries regarding the first scenario were easily achievable in all three languages (s. Listing 15).
@@
let $roleClasses := //RoleClass[@RefBaseClassPath=’PlantEquipment’]for $ie in //InternalElement where some $rc in $roleClasses satisfies $ie/RoleRequirements/@RefBaseRoleClassPath = string-join($rc/ancestor-or-self::*/@Name,’/’) return $ieSPARQL:
SELECT ?ieWHERE { ?ie rdf:type aml:InternalElement . ?ie aml:hasRoleRequirement/(rdfs:subClassOf)* aml:PlantEquipment .}
Both xPath and xQuery operate directly on the XML-tree, and have therefore no mechanisms to handle ReferencePaths. xQuery offers some mechanisms to circumvent this limitation: Technically, the second scenario could still be solved using xQuery. The approach includes reconstructing the reference path of all RoleClasses, iterating over them, and individually comparing them to the desired value (s. Listing 16). However, this approach is very limited in scope, e.g. capturing the open-ended transitive closure statement from Query 3 in this way is not possible. Using SPARQL, due to the support of Property Paths, all three queries can be answered in a much more compact and arguably more intuitive form (s. Listing 16).
@@
A SPARQL query that solves the second scenario can be found in Listing 18.
construct{?s aml:flows ?o.} where { ?s a aml:InternalElement. ?o a aml:InternalElement. ?s aml:hasInterface ?i1. ?o aml:hasInterface ?i2. ?i1 aml:isLinkedTo ?i2. OPTIONAL { ?s aml:hasAttribute ?vs. ?vs aml:hasRefAttributeType aml:ValveState. ?vs aml:hasAttributeValue ?state. } Filter(!bound(?state) || ?state = ’true’).}
@@ 7.2 Model Validation
In this use case, the validation concerns the correct usage of elements that were imported from user defined libraries. These libraries are commonly accompanied by textual specifications that define the proper usage of the imported elements. As those specifications lack a formal or machine interpretable definition, their correct usage needs to be manually checked by domain experts. For this use case, the AutomationML Application Recommendation Automation Project Configuration (AR APC) was considered. It is a framework that describes guidelines for the configuration of automation projects.[34]
@@
<rule context="caex:InternalElement[ caex:RoleRequirements[@RefBaseRoleClassPath= ’ArApcRcLib/SensorPort’]]"> <assert test=" every $link in caex:InternalLink satisfies (//caex:ExternalInterface[ @ID=$link/@RefPartnerSideA]/@RefBaseClassPath= ’ArApcIcLib/SensorPortInterface’ and //caex:ExternalInterface[ @ID=$link/@RefPartnerSideB]/@RefBaseClassPath= ’ArApcIcLib/SensorPortInterface’)"> </assert></rule>SHACL:
:SPIConnectionShape a sh:NodeShape ; sh:targetClass aml:SensorPortInterface; sh:property [ sh:path aml:isLinkedTo; sh:class aml:SensorPortInterface;] .